OpenStudy (anonymous):

Find the volume of a sphere with surface area equal to 100π ft^2.

OpenStudy (anonymous):

@goformit100 help me?

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

\[V=\frac{4}{3}\pi r^3\] \[A=4\pi r^2\] The surface area if \(100\pi\), so \[100\pi=4\pi r^2~~\iff~~25=r^2~~\iff~~r=?\] Find \(r\) and plug it into the volume formula.